i. Let A= [ 5 0 0]
[ 1 5 0]
[ 0 1 5]
Find a column vector X for which A= cX, for some c belong to R.

ii. Give an example with justification, of a skew-hermitian operator on C^2
1
Expert's answer
2021-07-20T09:22:05-0400

"A=\begin{vmatrix}
 5&0&0 \\
 1&5&0 \\0&1&5
\end{vmatrix}"

you cannot find the column X such that A=cX for c"\in" R

since A is 3x3 and cX is 3x1

however, we can find X such that AX=cX for some c"\in" A

clearly, 5 is an eign calue & A and the compounding eigen vector is

"X=\begin{bmatrix}
 0\\0\\a 
\end{bmatrix}"

foe any a"\in" R

we can check that AX=5X

"AX=\begin{bmatrix}
 5 & 0&0 \\1&5&0\\0&1&5
 \end{bmatrix}" "\begin{bmatrix}
 0\\0\\a 
\end{bmatrix}" ="X=\begin{bmatrix}
 0\\0\\5a 
\end{bmatrix}" ="5\begin{bmatrix}
 0\\0\\a 
\end{bmatrix}" ="5X"

as,"A=\begin{pmatrix}
 -1^0 & 2+1^0 \\
 -2+1^0 & 0
\end{pmatrix}"

"A^T=\begin{bmatrix}
 -1^0&-2+1^0\\2+1^0&0
\end{bmatrix}"

"A^T conjugate =(A^T)^*=\begin{bmatrix}
 1^0&-2+1^0\\2-1^0&0
\end{bmatrix}" "=" "=\begin{bmatrix}
 -1^0&2+1^0\\-2-1^0&0
\end{bmatrix}"


"\therefore" conjugate transpose of A is -A

Therefpre A is skew-hermitian